OpenStudy (anonymous):

Jill has been watching the squirrels that live in her neighborhood. The number of squirrels changes each week. n f(n) 1 8 2 12 3 18 4 27 Which function best shows the relationship between n and f(n)?

OpenStudy (anonymous):

f(n) = 12(1.5)n f(n) = 12(1.5)n-1 f(n) = 8(1.5)n f(n) = 8(1.5)n-1

OpenStudy (mathmale):

I'd suggest that you analyze this data: 1 8 2 12 3 18 4 27 and determine how each new value of f(n) is obtained from the previous value. Is this an arithmetic or a geometric sequence? How do you know?

OpenStudy (mathmale):

what is the first term of the given sequence? Which letter of the alphabet do we often choose to represent the first term?

OpenStudy (anonymous):

Umm is there another way of explain it cuz I don't really get it?

OpenStudy (mathmale):

the topic here is geometric sequences. Have you read up on that, either online or in any printed course materials that you may have?

OpenStudy (anonymous):

I do but I don't really get it

OpenStudy (mathmale):

First term: 8. Second term: 12. Given that this is a geometric (not arithmetic) sequence, what is the common ratio, r? Note: the common ratio, r, is the quantity by which we multiply the first term to get the second term. What is r?

OpenStudy (anonymous):

4

OpenStudy (mathmale):

So, you're saying that r is 4. Is 8(4) = 12? Again, 8 is the first term and 12 is the second term. If this is not correct, try again to find r. solve: 8r=12.

OpenStudy (anonymous):

You can times anything by 8 to get 12

OpenStudy (anonymous):

Wait 1.5

OpenStudy (mathmale):

Not in my experience. You need r. Your know that 8r = 12. Divide both sides of this equation by 8 to isolate r.

OpenStudy (mathmale):

Once y ou have a and r, the formula for the geometric sequence is \[a _{n+1}=a*r ^{n}\] with n beginning with 1, 2, 3, and so on.

OpenStudy (anonymous):

Okay but what is A

OpenStudy (mathmale):

I asked y ou earlier to identify the letter of the alphabet that is most often chosen to represent the first term of a geometric series. In this case the first term is what? The letter used to identify this first term is ' a ' .

OpenStudy (anonymous):

Sorry Im not getting what your asking

OpenStudy (anonymous):

Wait its D I got it